OpenStudy (anonymous):

A lottery has one $2000 prize, two $1000 prizes, and ten $500 prizes. Two thousand tickets are sold at $5 each. Find the expected value if a person buys two tickets

OpenStudy (amistre64):

expeted value is the sum of xP(x) if i read it right

OpenStudy (anonymous):

I got an answer of -2.80 is that correct...

OpenStudy (amistre64):

would the probabilities be: 1/2000 2/1999 10/1997 or 2000 denoms for all of them?

OpenStudy (anonymous):

first of all buying two tickets is a red herring.

OpenStudy (anonymous):

easy way to do this is imagine you buy all the tickets. you spend \[5\times \$2000=\$10,000\]

OpenStudy (anonymous):

that should be \[\$5\times 2000=\$10,000\] you make back \[\$2000+2\times $1000+10\times $500=\$9000\] for a loss of $1000

OpenStudy (anonymous):

averaged over the 2000 tickets you lost $4.5 per ticket.

OpenStudy (anonymous):

so if you buy one ticket expected loss is -4.5, for two tickets expect to lose -9

OpenStudy (anonymous):

we can do this using probability if you like

OpenStudy (anonymous):

hello zarkon!

OpenStudy (anonymous):

amistre can you please help with the other parts of the problem when you are done

OpenStudy (anonymous):

@gfoster did you check the formula?

OpenStudy (anonymous):

I'm not get the answer of -9

OpenStudy (anonymous):

really? let me check my arithmetic.

OpenStudy (anonymous):

$2,000+$2,000+10*$500=$9,000 yes?

OpenStudy (anonymous):

First I got the answer -2.8 then I got -6 I'm not sure what I'm doing worng

OpenStudy (anonymous):

yes that is correct

OpenStudy (anonymous):

do you know already what the answer is?

OpenStudy (anonymous):

No I don't - wish I did.

OpenStudy (anonymous):

because i get an expected value of -4.5 for each ticket, unless i messed up my arithmetic. oh i thought you knew my answer was wrong. i think it is right

OpenStudy (anonymous):

ya sorry i meant to say @satellite. the formula is correct

OpenStudy (anonymous):

you buy all the tickets. you spend $5*2000=$10,000 yes? but you only get back $9,000 is prize money for a net loss of $1,000

OpenStudy (anonymous):

I just didn't get the same answer that is why I'm asking I could be wrong - I'm not very good at this :-)

OpenStudy (anonymous):

oh ok a net loss of $1,000 averaged over the tickets. there are 2000 tickets

OpenStudy (anonymous):

i messed up on the average. -1000/2000=-.5

OpenStudy (anonymous):

you lose .50 per ticket

OpenStudy (anonymous):

so expected value is -.5 if you buy two tickets expect to lose -1

OpenStudy (anonymous):

if you buy 3 tickets expect to lose 1.5 and if you buy ten tickets expect to lose $5 etc. is it -.50 per ticket

OpenStudy (anonymous):

okay thank you for all the help.....

OpenStudy (anonymous):

tw

OpenStudy (anonymous):

if it is still not clear let me know and we can do it slowly

OpenStudy (anonymous):

sattelite those answers didnt come out right

OpenStudy (anonymous):

really? what is the formula for Ph?